# The KSBA moduli space of stable log Calabi–Yau surfaces

### 23 Feb 2024 v1

@@ 8. KSBA moduli spaces of stable log Calabi–Yau surfaces
the polarized mirror family as in Definition 6.17. In what follows we refer to this family as the “double mirror” to $(Y,D,L)$. In this section we first investigate in Theorem 8.2 how the double mirror changes if $(P,\mathscr{P})$ is rescaled. We then use this result to show in Theorem 8.6 that, up to rescaling $(P,\mathscr{P})$, the restriction of the double mirror to a particular one-dimensional locus is a toric degeneration. Using this, we show in Theorem 8.10 that the general fiber of the double mirror is deformation equivalent to $(Y,D,L)$. Consequently, we prove in Theorem 8.11 that we obtain a surjective and finite map from the toric variety $S^{\mathrm{sec}}_{\mathcal{X}_{\mathscr{P}}}$ to the KSBA moduli space $\mathcal{M}_{(Y,D,L)}$.
@@ 8.1. Base change of open Kulikov degenerations
Note that $\pi_{k\mathscr{P}}:\mathcal{X}_{k\mathscr{P}}\rightarrow\Delta$ is not an open Kulikov degeneration in general, as the total space $\mathcal{X}_{k\mathscr{P}}$ is singular. Instead, $\pi_{k\mathscr{P}}:\mathcal{X}_{k\mathscr{P}}\rightarrow\Delta$ is log smooth for the divisorial log structure on $\mathcal{X}_{k\mathscr{P}}$ defined by the central fiber $\mathcal{X}_{k\mathscr{P},0}:=\pi_{k\mathscr{P}}^{-1}(0)$, and the divisorial log structure on $\Delta$ defined by the origin $0\in\Delta$. Moreover, the dual intersection complex of the central fiber $\mathcal{X}_{k\mathscr{P},0}$ is the scaled polyhedral decomposition $k\mathscr{P}$ on the scaled polytope $nP$. The following result will allow us to construct a $K$-degeneration which is a resolution of singularities of $\pi_{k\mathscr{P}}:\mathcal{X}_{k\mathscr{P}}\rightarrow\Delta$.
@@
It suffices to identify the theta bases and the structure constants of the mirror algebra of $\pi_{\widetilde{\mathscr{P}}}:\mathcal{X}_{\widetilde{\mathscr{P}}}\rightarrow\Delta$ with the generators and structure constants of the subalgebra of the mirror algebra of $\pi_{\mathscr{P}}:\mathcal{X}_{\mathscr{P}}\rightarrow\Delta$ consisting of elements of degree divisible by $k$. To identify the theta bases, one notes that the index set of the theta basis for the mirror algebra of $\pi_{\widetilde{\mathscr{P}}}:\mathcal{X}_{\widetilde{\mathscr{P}}}\rightarrow\Delta$ is $C(kP)_{\mathbb{Z}}$, which is exactly the set of integral points of height divisible by $k$ is the cone $C(P)_{\mathbb{Z}}$ indexing the theta basis for the mirror algebra of $\pi_{\mathscr{P}}:\mathcal{X}_{\mathscr{P}}\rightarrow\Delta$. Hence, it remains to show that, for every $p,q,r\in C(kP)_{\mathbb{Z}}$ and $\beta\in NE(\mathcal{X}_{\mathscr{P}}/\overline{\mathcal{X}}_{\mathscr{P}})$, we have
@@
To prove Equation (8.1), we first remark that, even if the base change $\mathcal{X}_{k\mathscr{P}}$ of $\mathcal{X}_{\mathscr{P}}$ is not the total space of an open Kulikov degeneration, it is still log smooth over $\Delta$, and so one can also define invariants $N_{pqr}^{\beta}(\mathcal{X}_{k\mathscr{P}})$ for $\mathcal{X}_{k\mathscr{P}}$. As the base change $t\mapsto t^{k}$ does not change the central fiber and only scales the log structure, it follows that we have
@@
for all $p,q,r\in C(kP)_{\mathbb{Z}}$ and $\beta\in NE(\mathcal{X}_{k\mathscr{P}}/\overline{\mathcal{X}}_{k\mathscr{P}})\simeq NE(\mathcal{X}_{\mathscr{P}}/\overline{\mathcal{X}}_{\mathscr{P}})$. On the other hand, as $\nu:\mathcal{X}_{\widetilde{\mathscr{P}}}\rightarrow\mathcal{X}_{k\mathscr{P}}$ is a log modification, it follows from [55, §10] that
@@ 8.2. From semistable mirrors to toric degenerations
We use the notation introduced in the construction of $(P,\mathscr{P})$ in §2.2. For every $1\leq i\leq n$ and $1\leq j\leq N_{i}$, consider the image $L_{ij}$ in $P$ of the line in $\overline{P}$ passing through $p_{ij}$, and parallel to the edge $\overline{P}_{i}$ of $\overline{P}$. By Lemma 2.11, $L_{ij}$ is a non-trivial line segment in $P$, and so one can fix a rational point $p_{ij}^{\prime}\in L_{ij}$ close enough to $p_{ij}$ in the sense that the smallest cell of $\mathscr{P}$ containing $p_{ij}^{\prime}$ also contains $p_{ij}$, and such that all the points $p_{ij}^{\prime}$.
@@
Continuously moving the tips of the cut triangles $\Delta_{ij}$ on the line $L_{ij}$ away from $p_{ij}$ towards $p_{ij}^{\prime}$ produces a deformation of the integral affine structure on $P$, with endpoint an integral affine structure with distinct focus-focus singularities $\widetilde{p}_{ij}$ – see Figure 8.1. Moreover, one can assume that $\widetilde{p}_{ij}$ is contained in the interior of the edge of $\widetilde{\mathscr{P}}$ containing $p_{ij}$ and contained in $[p_{ij},p_{ij}^{\prime}]$. In particular, the cuts induced by the triangles $\Delta_{ij}$ deform into lines $\delta_{ij}$ connecting the side $\overline{P}_{i}$ to the focus-focus singularities $\widetilde{p}_{ij}$. Every integral point $v\in P_{\mathbb{Z}}$ is smooth for the deformed integral affine structure, and so its star defines a toric fan $\Sigma_{v}$. Note that, for every $i,j$ such that $v=p_{ij}$, the line segment $[p_{ij},p_{ij}^{\prime}]$ corresponds to a ray $\rho_{ij}$ of the fan $\Sigma_{v}$.
@@
Recall from §4.2 that the log Calabi–Yau surface $(X^{v}_{\widetilde{\mathscr{P}}},\partial X^{v}_{\widetilde{\mathscr{P}}})$ is a smoothing of the toric surface $(X^{v}_{n\mathscr{P}},\partial X^{v}_{n\mathscr{P}})$ with fan the star of $v$ in $\overline{P}$. It follows from the almost toric description of smoothing of quotient singularities – see [90, §6.1] and [32, Remark 2.7], that $(X^{v}_{\widetilde{\mathscr{P}}},\partial X^{v}_{\widetilde{\mathscr{P}}})$ admits an almost toric fibration over a Symington polytope obtained from a toric polytope of $(X^{v}_{k\overline{\mathscr{P}}},\partial X^{v}_{k\overline{\mathscr{P}}})$ by node smoothing. It follows from the combinatorial duality between internal blow-ups and node smoothings – see e.g.  [30, p 488], that the pair $(X^{v}_{\widetilde{\mathscr{P}}},\partial X^{v}_{\widetilde{\mathscr{P}}})$ is diffeomorphic, and so deformation-equivalent by [34, Theorem 5.14], to the log Calabi–Yau surface described in Lemma 8.3 as internal blow-ups of the toric surface $(X_{\Sigma_{v}},\partial X_{\Sigma_{v}})$. ∎
@@
Since $\widetilde{\mathscr{P}}$ is a good polyhedral decomposition on $kP$ as in §2.3, there is a strictly convex integral multi-valued PL function $\varphi$ on $kP$ whose domains of linearity agree with $\widetilde{\mathscr{P}}$, as in §2.4. Then, we obtain from $\varphi$ a strictly convex integral multi-valued PL function $\varphi^{gs}$ on $kP$ on the integral affine manifold with singularities $(kP,\widetilde{\mathscr{P}})^{gs}$ obtained from $(kP,\widetilde{\mathscr{P}})$ by pushing singularities away from the vertices as described before Lemma 8.3 – see [18, §3] where such PL functions on integral affine manifolds with singularities pushed away are studied in detail. Therefore, by [46, Theorem 2.34] since $P$ is a disk with second cohomology zero, $\mathcal{X}_{\widetilde{\mathscr{P}},0}^{gs}$ is a quasi-projective scheme. Hence, we have an ample line bundle $L_{0}$ on $\mathcal{X}_{\widetilde{\mathscr{P}},0}^{gs}$ whose pullback under $\pi_{\widetilde{\mathscr{P}}}$ is a nef line bundle on $\mathcal{X}_{\widetilde{\mathscr{P}}}$. Since $H^{2}(\mathcal{X}_{\widetilde{\mathscr{P}},0},\mathcal{O}_{\mathcal{X}_{\widetilde{\mathscr{P}},0}})=0$ by Lemma 3.10, this line bundle deforms to a nef line bundle $L$ on $\mathcal{X}_{\widetilde{\mathscr{P}}}$. Since $\mathcal{X}_{\widetilde{\mathscr{P}}}$ is a Mori dream space by Proposition 3.36, there is a projective birational morphism $p:\mathcal{X}_{\widetilde{\mathscr{P}}}\rightarrow\mathcal{X}_{\widetilde{\mathscr{P}}}^{gs}$ exactly contracting the curves $C$ with $L\cdot C=0$, that is the finitely many internal exceptional curves in the central fiber contracted by $p_{0}$.
@@ 8.3. Toric degenerations and intrinsic mirrors
By Lemma 8.4, for $k$ large enough, there exists a refinement $\widetilde{\mathscr{P}}$ of the scaled polyhedral decomposition $k\mathscr{P}$ such that the corresponding quasi-projective Kulikov degeneration $\mathcal{X}_{\widetilde{\mathscr{P}}}\rightarrow\Delta$ admits a birational contraction to a toric degeneration $\mathcal{X}_{\widetilde{\mathscr{P}}}^{gs}\rightarrow\Delta$ in the sense of [46]. A mirror construction for so-called positive and simple polarized toric degenerations of Calabi–Yau varieties has been introduced by Gross–Siebert [46] – see [38] for an exposition in dimension two. The toric degeneration $\mathcal{X}_{\widetilde{\mathscr{P}}}^{gs}\rightarrow\Delta$ is an example of positive and simple toric degeneration, since its dual intersection complex is an integral affine manifold with only focus-focus singularities localized in the interior of the edges of the polyhedral decomposition. Unlike the intrinsic mirror construction for log smooth degenerations [47, 48] reviewed in §5 and used in §6-7, which is based on the enumerative geometry of punctured Gromov–Witten invariants, the mirror construction for toric degenerations is based on a scattering diagram recursively constructed from explicit initial data, and so is mostly combinatorial. In particular, it is often easier to determine the output of the mirror construction for toric degenerations. In this section, we compare the intrinsic mirror family $(\mathcal{Y}_{\mathcal{X}_{\widetilde{\mathscr{P}}}},\mathcal{D}_{\mathcal{X}_{\widetilde{\mathscr{P}}}},\mathcal{L}_{\widetilde{\mathscr{P}}})$ to $\mathcal{X}_{\widetilde{\mathscr{P}}}\rightarrow\Delta$ with the mirror to the toric degeneration $\mathcal{X}_{\widetilde{\mathscr{P}}}^{gs}\rightarrow\Delta$. Similar comparison results have been obtained in [40, §3.3], [16], [70, §6.1], [17, §2].
Let $G$ be the face of $NE(\mathcal{X}_{\widetilde{\mathscr{P}}}/\overline{\mathcal{X}}_{\widetilde{\mathscr{P}}})$ spanned by the finitely many internal exceptional curves blow down by $p:\mathcal{X}_{\widetilde{\mathscr{P}}}\rightarrow\mathcal{X}_{\widetilde{\mathscr{P}}}^{gs}$ as in Lemmas 8.3 and 8.4. Consider the monoid ideal $J_{G}=NE(\mathcal{X}_{\widetilde{\mathscr{P}}}/\overline{\mathcal{X}}_{\widetilde{\mathscr{P}}})\setminus G$, and similarly as in [40, Definition 3.14] define the Gross–Siebert locus as the open torus $T^{gs}=\mathrm{Spec}\,\mathbb{C}[G^{\mathrm{gp}}]$ in the toric stratum $\mathrm{Spec}\,\mathbb{C}[G]=\mathrm{Spec}\,\mathbb{C}[NE(\mathcal{X}_{\widetilde{\mathscr{P}}}/\overline{\mathcal{X}}_{\widetilde{\mathscr{P}}})/J_{G}]$ of the toric variety $S_{\mathcal{X}_{\widetilde{\mathscr{P}}}}$.
@@
On the other hand, the choice of $H$ naturally defines a strictly convex multivalued PL function $\varphi^{gs}_{H}$ on the integral affine manifold with singularities $(kP,\widetilde{\mathscr{P}})^{gs}$ obtained from $(kP,\widetilde{\mathscr{P}})$ by generically pushing away the singularities from the vertices to irrational points on edges of $\widetilde{\mathscr{P}}$ as described before Lemma 8.3. Indeed, by standard toric geometry, the ample divisor $H$ restricted to each of the toric irreducible components of $\mathcal{X}_{\widetilde{\mathscr{P}},0}^{gs}$ defines strictly convex PL functions around the various vertices of $\widetilde{\mathscr{P}}$. Moreover, the choice of the point $a\in T^{gs}$, viewed as an algebra homomorphism $a:\mathbb{C}[G^{gp}]\rightarrow\mathbb{C}$, naturally defines functions $1+a(t^{E_{i}})z^{\pm m_{i}}$ attached to the monodromy invariant directions ±miplus-or-$\pm m_{i}$ of the focus-focus singularity associated to the contracted curve $E_{i}$. Taking as input the integral affine $(kP,\widetilde{\mathscr{P}})^{gs}$, having only focus-focus singularities placed at irrational points on the edges of $\widetilde{\mathscr{P}}$, the strictly convex multivalued PL function $\varphi^{gs}_{H}$, and the functions $1+a(t^{E_{i}})z^{\pm m_{i}}$ as “slab functions”, the mirror construction for toric degeneration of [46] – see also [38, Chapter 6], produces a projective family $\hat{\mathcal{Y}}^{gs}_{H,a}\rightarrow\mathrm{Spec}\,\mathbb{C}[\![t]\!]$, with a divisor $\hat{\mathcal{D}}_{H,a}^{gs}$, and an ample line bundle ^$\hat{\mathcal{L}}^{gs}_{H,a}$.
@@
For every ample divisor $H$ on $\mathcal{X}^{gs}_{\widetilde{\mathscr{P}}}$, and every point $a\in T^{gs}$, the mirror family $(\hat{\mathcal{Y}}^{gs}_{H,a},\hat{\mathcal{D}}_{H,a}^{gs},\hat{\mathcal{L}}^{gs}_{H,a})\rightarrow\mathrm{Spec}\,\mathbb{C}[\![t]\!]$ of the toric degeneration $\mathcal{X}^{gs}_{\widetilde{\mathscr{P}}}\rightarrow\Delta$ canonically extends to a family $(\mathcal{Y}^{gs}_{H,a},\mathcal{D}_{H,a}^{gs},\mathcal{L}^{gs}_{H,a})\rightarrow\mathbb{A}^{1}$, which is isomorphic to the pullback, along the map
$(\mathbb{A}^{1})_{H,a}\longrightarrow S_{\mathcal{X}_{\widetilde{\mathscr{P}}}}$
in Equation (8.4), of the mirror family $(\mathcal{Y}_{\mathcal{X}_{\widetilde{\mathscr{P}}}},\mathcal{D}_{\mathcal{X}_{\widetilde{\mathscr{P}}}},\mathcal{L}_{\mathcal{X}_{\widetilde{\mathscr{P}}}})\rightarrow S_{\mathcal{X}_{\widetilde{\mathscr{P}}}}$ of the quasi-projective open Kulikov degeneration $\mathcal{X}_{\widetilde{\mathscr{P}}}\rightarrow\Delta$.

### 30 Mar 2024 v2

@@ 8. KSBA moduli spaces of stable log Calabi–Yau surfaces
the polarized mirror family as in Definition 6.17. In what follows we refer to this family as the “double mirror” to $(Y,D,L)$. In this section we first investigate in Theorem 8.2 how the double mirror changes if $(P,\mathscr{P})$ is rescaled. We then use this result to show in Theorem 8.6 that, up to rescaling $(P,\mathscr{P})$, the restriction of the double mirror to a particular one-dimensional locus is a toric degeneration. Using this, we show in Theorem 8.10 that the general fiber of the double mirror is deformation equivalent to $(Y,D,L)$. Consequently, we prove in Theorem 8.11 that we obtain a surjective and finite map from the toric variety $S^{\mathrm{sec}}_{\mathcal{X}_{\mathscr{P}}}$ to the coarse KSBA moduli space $M_{(Y,D,L)}$.
@@ 8.1. Base change of open Kulikov degenerations
Note that $\pi_{k\mathscr{P}}:\mathcal{X}_{k\mathscr{P}}\rightarrow\Delta$ is not an open Kulikov degeneration in general, as the total space $\mathcal{X}_{k\mathscr{P}}$ is singular. Instead, $\pi_{k\mathscr{P}}:\mathcal{X}_{k\mathscr{P}}\rightarrow\Delta$ is log smooth for the divisorial log structure on $\mathcal{X}_{k\mathscr{P}}$ defined by the central fiber $\mathcal{X}_{k\mathscr{P},0}:=\pi_{k\mathscr{P}}^{-1}(0)$, and the divisorial log structure on $\Delta$ defined by the origin $0\in\Delta$. Moreover, the dual intersection complex of the central fiber $\mathcal{X}_{k\mathscr{P},0}$ is the scaled polyhedral decomposition $k\mathscr{P}$ on the scaled polytope $kP$. The following result will allow us to construct an open Kulikov degeneration which is a resolution of singularities of $\pi_{k\mathscr{P}}:\mathcal{X}_{k\mathscr{P}}\rightarrow\Delta$.
@@
It suffices to identify the theta bases and the structure constants of the mirror algebra of $\pi_{\widetilde{\mathscr{P}}}:\mathcal{X}_{\widetilde{\mathscr{P}}}\rightarrow\Delta$ with the generators and structure constants of the subalgebra of the mirror algebra of $\pi_{\mathscr{P}}:\mathcal{X}_{\mathscr{P}}\rightarrow\Delta$ consisting of elements of degree divisible by $k$. To identify the theta bases, one notes that the index set of the theta basis for the mirror algebra of $\pi_{\widetilde{\mathscr{P}}}:\mathcal{X}_{\widetilde{\mathscr{P}}}\rightarrow\Delta$ is $C(kP)_{\mathbb{Z}}$, which is exactly the set of integral points of height divisible by $k$ in the cone $C(P)_{\mathbb{Z}}$ indexing the theta basis for the mirror algebra of $\pi_{\mathscr{P}}:\mathcal{X}_{\mathscr{P}}\rightarrow\Delta$. Hence, it remains to show that, for every $p,q,r\in C(kP)_{\mathbb{Z}}$ and $\beta\in NE(\mathcal{X}_{\mathscr{P}}/\overline{\mathcal{X}}_{\mathscr{P}})$, we have
@@
To prove Equation (8.1), we first note that, even if the base change $\mathcal{X}_{k\mathscr{P}}$ of $\mathcal{X}_{\mathscr{P}}$ is not the total space of an open Kulikov degeneration, it is still log smooth over $\Delta$, and so one can also define invariants $N_{pqr}^{\beta}(\mathcal{X}_{k\mathscr{P}})$ for $\mathcal{X}_{k\mathscr{P}}$. As the base change $t\mapsto t^{k}$ does not change the central fiber and only scales the log structure, it follows that we have
@@
for all $p,q,r\in C(kP)_{\mathbb{Z}}$ and $\beta\in NE(\mathcal{X}_{k\mathscr{P}}/\overline{\mathcal{X}}_{k\mathscr{P}})\simeq NE(\mathcal{X}_{\mathscr{P}}/\overline{\mathcal{X}}_{\mathscr{P}})$. On the other hand, as $\nu:\mathcal{X}_{\widetilde{\mathscr{P}}}\rightarrow\mathcal{X}_{k\mathscr{P}}$ is a log modification, it follows from [56, §10] that
@@ 8.2. From semistable mirrors to toric degenerations
We use the notation introduced in the construction of $(P,\mathscr{P})$ in §2.2. For every $1\leq i\leq n$ and $1\leq j\leq N_{i}$, consider the image $L_{ij}$ in $P$ of the line in $\overline{P}$ passing through $p_{ij}$, and parallel to the edge $\overline{P}_{i}$ of $\overline{P}$. By Lemma 2.11, the intersection of $L_{ij}$ with the interior of $P$ is an interval of positive length, and so one can find a rational point $p_{ij}^{\prime}\in L_{ij}$ close enough to $p_{ij}$ in the sense that the smallest cell of $\mathscr{P}$ containing $p_{ij}^{\prime}$ also contains $p_{ij}$, and such that all the points $p_{ij}^{\prime}$ are distinct and contained in the interior of $P$.
@@
Continuously moving the tips of the cut triangles $\Delta_{ij}$ on the line $L_{ij}$ away from $p_{ij}$ towards $p_{ij}^{\prime}$ produces a deformation of the integral affine structure on $P$, with endpoint an integral affine structure with distinct focus-focus singularities $\widetilde{p}_{ij}$ – see Figure 8.1. Moreover, we can assume that $\widetilde{p}_{ij}$ is contained in the interior of the edge of $\widetilde{\mathscr{P}}$ containing $p_{ij}$ and contained in $[p_{ij},p_{ij}^{\prime}]$. In particular, the cuts induced by the triangles $\Delta_{ij}$ deform into lines $\delta_{ij}$ connecting the side $\overline{P}_{i}$ to the focus-focus singularities $\widetilde{p}_{ij}$. Every integral point $v\in P_{\mathbb{Z}}$ is smooth for the deformed integral affine structure, and so its star defines a toric fan $\Sigma_{v}$. Note that, for every $i,j$ such that $v=p_{ij}$, the line segment $[p_{ij},p_{ij}^{\prime}]$ corresponds to a ray $\rho_{ij}$ of the fan $\Sigma_{v}$.
@@
Recall from §4.2 that the log Calabi–Yau surface $(X^{v}_{\widetilde{\mathscr{P}}},\partial X^{v}_{\widetilde{\mathscr{P}}})$ is a smoothing of the toric surface $(X^{v}_{n\mathscr{P}},\partial X^{v}_{n\mathscr{P}})$ with fan the star of $v$ in $\overline{P}$. Using the almost toric description of smoothings of quotient singularities – see [91, §6.1] and [32, Remark 2.7], it follows that $(X^{v}_{\widetilde{\mathscr{P}}},\partial X^{v}_{\widetilde{\mathscr{P}}})$ admits an almost toric fibration over a Symington polytope obtained from a toric polytope of $(X^{v}_{k\overline{\mathscr{P}}},\partial X^{v}_{k\overline{\mathscr{P}}})$ by node smoothing. Therefore, by the combinatorial duality between internal blow-ups and node smoothings – see e.g.  [30, p 488], the pair $(X^{v}_{\widetilde{\mathscr{P}}},\partial X^{v}_{\widetilde{\mathscr{P}}})$ is diffeomorphic, and so deformation-equivalent by [34, Theorem 5.14], to the log Calabi–Yau surface described in Lemma 8.3 as internal blow-ups of the toric surface $(X_{\Sigma_{v}},\partial X_{\Sigma_{v}})$. ∎
@@
Since $\widetilde{\mathscr{P}}$ is a good polyhedral decomposition on $kP$ as in §2.3, there is a strictly convex integral multi-valued PL function $\varphi$ on $kP$ whose domains of linearity agree with $\widetilde{\mathscr{P}}$, as in §2.4. Then, we obtain from $\varphi$ a strictly convex integral multi-valued PL function $\varphi^{gs}$ on $kP$ on the integral affine manifold with singularities $(kP,\widetilde{\mathscr{P}})^{gs}$ obtained from $(kP,\widetilde{\mathscr{P}})$ by pushing singularities away from the vertices as described before Lemma 8.3 – see [18, §3] where such PL functions on integral affine manifolds with singularities pushed away are studied in detail. Therefore, by [47, Theorem 2.34] since $P$ is a disk with second cohomology zero, $\mathcal{X}_{\widetilde{\mathscr{P}},0}^{gs}$ is a quasi-projective scheme. Hence, we have an ample line bundle $L_{0}$ on $\mathcal{X}_{\widetilde{\mathscr{P}},0}^{gs}$ whose pullback under $\pi_{\widetilde{\mathscr{P}}}$ is a nef line bundle on $\mathcal{X}_{\widetilde{\mathscr{P}}}$. Since $H^{2}(\mathcal{X}_{\widetilde{\mathscr{P}},0},\mathcal{O}_{\mathcal{X}_{\widetilde{\mathscr{P}},0}})=0$ by Lemma 3.10, this line bundle deforms to a nef line bundle $L$ on $\mathcal{X}_{\widetilde{\mathscr{P}}}$. Since $\mathcal{X}_{\widetilde{\mathscr{P}}}$ is a Mori dream space by Proposition 3.36, there is a projective birational morphism $p:\mathcal{X}_{\widetilde{\mathscr{P}}}\rightarrow\mathcal{X}_{\widetilde{\mathscr{P}}}^{gs}$ exactly contracting the curves $C$ with $L\cdot C=0$, that is the finitely many internal exceptional curves in the central fiber contracted by $p_{0}$.
@@ 8.3. Toric degenerations and intrinsic mirrors
By Lemma 8.4, for $k$ large enough, there exists a refinement $\widetilde{\mathscr{P}}$ of the scaled polyhedral decomposition $k\mathscr{P}$ such that the corresponding quasi-projective Kulikov degeneration $\mathcal{X}_{\widetilde{\mathscr{P}}}\rightarrow\Delta$ admits a birational contraction to a toric degeneration $\mathcal{X}_{\widetilde{\mathscr{P}}}^{gs}\rightarrow\Delta$ in the sense of [47]. A mirror construction for so-called positive and simple polarized toric degenerations of Calabi–Yau varieties has been introduced by Gross–Siebert [47] – see [39] for an exposition in dimension two. The toric degeneration $\mathcal{X}_{\widetilde{\mathscr{P}}}^{gs}\rightarrow\Delta$ is an example of positive and simple toric degeneration, since its dual intersection complex is an integral affine manifold with only focus-focus singularities localized in the interior of the edges of the polyhedral decomposition. Unlike the intrinsic mirror construction for log smooth degenerations [48, 49] reviewed in §5 and used in §6-7, which is based on the enumerative geometry of punctured Gromov–Witten invariants, the mirror construction for toric degenerations is based on a scattering diagram recursively constructed from explicit initial data, and so is mostly combinatorial. In particular, it is often easier to determine the output of the mirror construction for toric degenerations. In this section, we compare the intrinsic mirror family $(\mathcal{Y}_{\mathcal{X}_{\widetilde{\mathscr{P}}}},\mathcal{D}_{\mathcal{X}_{\widetilde{\mathscr{P}}}},\mathcal{L}_{\widetilde{\mathscr{P}}})$ to $\mathcal{X}_{\widetilde{\mathscr{P}}}\rightarrow\Delta$ with the mirror to the toric degeneration $\mathcal{X}_{\widetilde{\mathscr{P}}}^{gs}\rightarrow\Delta$. Similar comparison results have been obtained in [41, §3.3], [16], [71, §6.1], [17, §2].
Let $G$ be the face of $NE(\mathcal{X}_{\widetilde{\mathscr{P}}}/\overline{\mathcal{X}}_{\widetilde{\mathscr{P}}})$ spanned by the finitely many internal exceptional curves blow down by $p:\mathcal{X}_{\widetilde{\mathscr{P}}}\rightarrow\mathcal{X}_{\widetilde{\mathscr{P}}}^{gs}$ as in Lemmas 8.3 and 8.4. Consider the monoid ideal $J_{G}=NE(\mathcal{X}_{\widetilde{\mathscr{P}}}/\overline{\mathcal{X}}_{\widetilde{\mathscr{P}}})\setminus G$, and similarly as in [41, Definition 3.14] define the Gross–Siebert locus as the open torus $T^{gs}=\mathrm{Spec}\,\mathbb{C}[G^{\mathrm{gp}}]$ in the toric stratum $\mathrm{Spec}\,\mathbb{C}[G]=\mathrm{Spec}\,\mathbb{C}[NE(\mathcal{X}_{\widetilde{\mathscr{P}}}/\overline{\mathcal{X}}_{\widetilde{\mathscr{P}}})/J_{G}]$ of the toric variety $S_{\mathcal{X}_{\widetilde{\mathscr{P}}}}$.
@@
On the other hand, the choice of $H$ naturally defines a strictly convex multivalued PL function $\varphi^{gs}_{H}$ on the integral affine manifold with singularities $(kP,\widetilde{\mathscr{P}})^{gs}$ obtained from $(kP,\widetilde{\mathscr{P}})$ by generically pushing away the singularities from the vertices to irrational points on edges of $\widetilde{\mathscr{P}}$ as described before Lemma 8.3. Indeed, by standard toric geometry, the ample divisor $H$ restricted to each of the toric irreducible components of $\mathcal{X}_{\widetilde{\mathscr{P}},0}^{gs}$ defines strictly convex PL functions around the various vertices of $\widetilde{\mathscr{P}}$. Moreover, the choice of the point $a\in T^{gs}$, viewed as an algebra homomorphism $a:\mathbb{C}[G^{gp}]\rightarrow\mathbb{C}$, naturally defines functions $1+a(t^{E_{i}})z^{\pm m_{i}}$ attached to the monodromy invariant directions ±miplus-or-$\pm m_{i}$ of the focus-focus singularity associated to the contracted curve $E_{i}$. Taking as input the integral affine $(kP,\widetilde{\mathscr{P}})^{gs}$, having only focus-focus singularities placed at irrational points on the edges of $\widetilde{\mathscr{P}}$, the strictly convex multivalued PL function $\varphi^{gs}_{H}$, and the functions $1+a(t^{E_{i}})z^{\pm m_{i}}$ as “slab functions”, the mirror construction for toric degeneration of [47] – see also [39, Chapter 6], produces a projective family $\hat{\mathcal{Y}}^{gs}_{H,a}\rightarrow\mathrm{Spec}\,\mathbb{C}[\![t]\!]$, with a divisor $\hat{\mathcal{D}}_{H,a}^{gs}$, and an ample line bundle ^$\hat{\mathcal{L}}^{gs}_{H,a}$.
@@
For every ample divisor $H$ on $\mathcal{X}^{gs}_{\widetilde{\mathscr{P}}}$, and every point $a\in T^{gs}$, the mirror family $(\hat{\mathcal{Y}}^{gs}_{H,a},\hat{\mathcal{D}}_{H,a}^{gs},\hat{\mathcal{L}}^{gs}_{H,a})\rightarrow\mathrm{Spec}\,\mathbb{C}[\![t]\!]$ of the toric degeneration $\mathcal{X}^{gs}_{\widetilde{\mathscr{P}}}\rightarrow\Delta$ canonically extends to a family $(\mathcal{Y}^{gs}_{H,a},\mathcal{D}_{H,a}^{gs},\mathcal{L}^{gs}_{H,a})\rightarrow\mathbb{A}^{1}$, which is isomorphic to the pullback, along the map $(\mathbb{A}^{1})_{H,a}\longrightarrow S_{\mathcal{X}_{\widetilde{\mathscr{P}}}}$ in (8.4), of the mirror family $(\mathcal{Y}_{\mathcal{X}_{\widetilde{\mathscr{P}}}},\mathcal{D}_{\mathcal{X}_{\widetilde{\mathscr{P}}}},\mathcal{L}_{\mathcal{X}_{\widetilde{\mathscr{P}}}})\rightarrow S_{\mathcal{X}_{\widetilde{\mathscr{P}}}}$ of the quasi-projective open Kulikov degeneration $\mathcal{X}_{\widetilde{\mathscr{P}}}\rightarrow\Delta$.